OpenStudy (anonymous):

Find the derivative of the function using the definition of derivative. the square root of 8-x

OpenStudy (dumbcow):

oh the dreaded limit definition of derivatives

OpenStudy (dumbcow):

haha can i just teach you the shortcut lol

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

no that wont help you on a test

OpenStudy (anonymous):

I know that you pull down the number that is above the exponent so the x would disappear and become 1 but would the 8 also disappear?

OpenStudy (dumbcow):

so you have \[\lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{h} = \lim_{h \rightarrow 0} \frac{\sqrt{8-(x+h)} - \sqrt{8-x}}{h}\]

OpenStudy (dumbcow):

ahh yes the 8 would disappear .... derivative of constant is 0

OpenStudy (anonymous):

so it would just be the square root of 1?

OpenStudy (dumbcow):

no because the whole thing is inside of sqare root, its a bit more complicated

OpenStudy (anonymous):

Now im confused

OpenStudy (dumbcow):

anyway are you talking about limit definition here?

OpenStudy (dumbcow):

or where are you in your calculus course? this way i know the correct approach to help you

OpenStudy (anonymous):

I am on chapter 2.8

OpenStudy (dumbcow):

whats the chapter heading ...

OpenStudy (anonymous):

Nevermind i figured it out. Thank you

OpenStudy (dumbcow):

alrighty then